# Thresholded Lasso for high dimensional variable selection

### v2

@@ 1 Introduction
Here $X$ is an $n\times p$ design matrix, $Y$ is a vector of noisy observations, and $\epsilon$ is the noise term. We assume throughout this paper that $p\geq n$ (i.e. high-dimensional) and $\epsilon$ is a vector of i.i.d. normal $N(0,\sigma^{2})$ random variables. Denote by $[p]=\{1,\ldots,p\}$. The notation $\left\lVert\beta\right\rVert_{2}=(\sum_{j}\beta_{j}^{2})^{1/2}$ stands for the $\ell_{2}$ norm of $\beta$. Given such a linear model, two key tasks are: (1) to select the relevant set of variables and (2) to estimate $\beta$ with bounded $\ell_{2}$ loss. In particular, recovery of the sparsity pattern $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta):=\left\{j\,:\,\beta_{j}\neq 0\right\}$, also known as variable (model) selection, refers to the task of correctly identifying the support set, or a subset of “significant” coefficients in $\beta$, based on the noisy observations. Here and in the sequel, we assume that each column vector $X_{j}\in\mathbb{R}^{n},j\in[p]$ of the fixed design matrix $X$ has $\ell_{2}$-length $\left\lVert X_{j}\right\rVert_{2}=\sqrt{n}$.
[…]
@@
[…]
@@
[…]
@@
In Theorem 2.1, we show that the critical threshold level for estimating a high dimensional sparse vector $\beta\in\mathbb{R}^{p}$ should be set at the level $t_{0}\asymp\lambda\sigma$ for $\lambda=\sqrt{2\log p/n}$, to retain signals in $\beta_{\text{\rm init}}$ at or above the level of $\lambda\sigma$, where $\beta_{\text{\rm init}}$ is the solution to the Lasso estimator (2) obtained in Step 1. Moreover, we show that $d_{0}$ and $t_{0}$ are allowed to depend on sparse and restricted eigenvalue parameters of the design matrix; cf. Section 3.1. From a theoretical point of view, the framework for our analysis is set upon the Restricted Eigenvalue type of condition and an upper sparse eigenvalue condition, namely,
@@
$\Lambda_{\max}(2s)\;\stackrel{{\scriptstyle\triangle}}{{=}}\;\max_{\upsilon\not=0;2s-\text{sparse}}\;\left\lVert X\upsilon\right\rVert_{2}^{2}/(n\left\lVert\upsilon\right\rVert_{2}^{2})<\infty.$ (5)
@@
These are among the most general assumptions on the design matrix, guaranteeing sparse recovery and oracle inequalities in the $\ell_{2}$ loss for the Lasso estimator as well as the Dantzig selector: $\textnormal{{RE}}(s,k_{0},X)$ is shown to be a relaxation of the restricted isometry property (RIP) under suitable choices of parameters involved in each condition [9, 3].
@@
Part of this work was presented in a conference paper by [47]. Importantly, we present significant and novel extensions in theory and numerical simulations, with regards to the Thresholded Lasso and the Lasso under the RE and sparse eigenvalue conditions. While the crucial theoretical and methodological ideas presented here originate from [47], the current work significantly expands the original ideas and show new results on the sparse oracle inequalities in Theorems 2.1 and 2.4. Compared to the original paper [47], we further study the behavior of the Thresholded Lasso in several challenging situations in Sections 45, and the supplementary Section K, and show that our estimator is robust and adaptive to the overwhelming presence of weak signals in $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ in both theoretical and practical senses. We show that the Thresholded Lasso tradeoffs false positives and false negatives nicely in this case: its advantage in terms of model selection over the Lasso and adaptive Lasso [52, 51] is clearly evident by examining their ROC curves empirically. Our numerical simulations in Section K show that the rates for exact recovery of the support rise sharply for a few types of random matrices once the number of samples passes a certain threshold, using the Thresholded Lasso estimator.
[…]
@@
$\Lambda_{\min}(2s)\;\stackrel{{\scriptstyle\triangle}}{{=}}\;\min_{\upsilon\not=0;2s-\text{sparse}}\;\;\left\lVert X\upsilon\right\rVert_{2}^{2}/(n\left\lVert\upsilon\right\rVert_{2}^{2})>0,$ (6)
@@
${\left\lvert{\;\langle{\,X_{T}\upsilon,X_{T^{\prime}}{\upsilon^{\prime}}\,}\rangle\;}\right\rvert}/{n}\leq\theta_{s,s^{\prime}}\left\lVert\upsilon\right\rVert_{2}\left\lVert\upsilon^{\prime}\right\rVert_{2},\;\text{ where }s+s^{\prime}\leq p.$ (7)
[…]
@@ 1.1 Sparse oracle inequalities
Specifically, (a) we wish to obtain $\hat{\beta}$ such that $|\mathop{\text{\rm supp}\kern 0.86108pt}(\hat{\beta})\setminus S|$ (and sometimes the set difference between $S$ and $\mathop{\text{\rm supp}\kern 0.86108pt}(\hat{\beta})$ denoted by $|S\triangle\mathop{\text{\rm supp}\kern 0.86108pt}(\hat{\beta})|$ also) is small, with high probability; (b) while at the same time, we wish to bound $\lVert\hat{\beta}-\beta\rVert_{2}^{2}$, within logarithmic factor of the ideal mean squared error one would achieve with an oracle that would supply perfect information about which coordinates are non-zero and which are above the noise level (hence achieving the oracle inequality as studied by [13] and [9]). Here we denote by $\beta_{I}$ the restriction of $\beta$ to the set $I$.
Formally, we evaluate the selection set through the following criterion. Consider the least squares estimators $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, where $I\subset[p]$ and $|I|\ \leq s$. Here and in the sequel, let $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)=\hat{\beta}_{I}$ and $\hat{\beta}_{I^{c}}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)=0$. Consider the ideal least-squares estimator $\beta^{\diamond}$ based on a subset $I$ of size at most $s$, which minimizes the mean squared error:
@@
${\textbf{E}}\left\lVert\beta-\beta^{\diamond}\right\rVert_{2}^{2}$$\geq$ $\min\left(1,{1}/{\Lambda_{\max}(s)}\right)\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n),$ (10)
where $\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n)=\min_{I\subseteq[p]}\left\lVert\beta-\beta_{I}\right\rVert_{2}^{2}+{|I|\sigma^{2}}/{n}$
@@
in view of  (10); cf. Remark 2.2 and the supplementary Section D. Here the $\ell_{2}$ loss in (11) is optimal up to a $\log p$ factor. We note that (12) is not the tightest upper bound that we could derive due to a relaxation we have on the lower bound as stated in (10). Nevertheless, we use it for its simplicity.
Essential sparsity and objectives. The current paper answers the following question: Is there a good thresholding rule that enables us to obtain a sufficiently sparse estimator $\hat{\beta}$ that satisfies an oracle inequality in the sense of (11), when some components of $\beta_{S}$ are well below $\sigma/\sqrt{n}$? Such oracle results are accomplished without any knowledge of the significant coordinates or parameter values of $\beta$. Both Theorem 2.1 and the supplementary Theorem I.3 answer this question positively, where we elaborate upon the sparse recovery properties of the Lasso and Dantzig selector in combination with thresholding and refitting.
For a given pair of $(n,p)$ values, the essential sparsity parameter $s_{0}$ characterizes more accurately than $s$ the number of significant coefficients of $\beta$ with respect to the noise level $\sigma$ that we should (could) try to recover. Denote by $s_{0}$ the smallest integer such that the following holds [9]:
@@
The parameter $s_{0}$ is relevant especially when we do not wish to impose any lower bound on $\beta_{\min}$. This is precisely the focus of Theorem 2.1. To make this statement precise, we have as a consequence of the definition in (13),
@@
The Thresholded Lasso algorithm is constructive in that it relies neither on the unknown parameters $\left\lvert S\right\rvert$ or $\beta_{\min}:=\min_{j\in S}|\beta_{j}|$, nor the exact knowledge of those that characterize the incoherence conditions on $X$. Instead, our choices of $\lambda_{n}$ and thresholding parameter $t_{0}$ only depend on $\sigma,n$, and $p$, and some crude estimation of certain sparse eigenvalue parameters; cf. Section 3.1. In practical settings, one can choose $\lambda_{n}$ using cross-validation; See for example the subsequent work by [50], where we use cross-validation to choose both penalty and threshold parameters in the context of covariance selection based on Gaussian graphical models.
In Section 4, we briefly discuss possibilities of recovering a subset of strong signals via thresholding, despite the existence of (or influence from) other relatively weaker signals. Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. As a consequence of the definition in (13), we have $T_{0}=\{1,\ldots,s_{0}\}$ and $|\beta_{j}|<\lambda\sigma$ for all $j\in T_{0}^{c}$ (cf. (14)). More precisely, we decompose $T_{0}=\{1,\ldots,s_{0}\}$ into two sets: $A_{0}$ and $T_{0}\setminus A_{0}$, where $A_{0}$ contains the set of coefficients of $\beta$ strictly above $\lambda\sigma$, for which we define a constant $\beta_{\min,A_{0}}$:
@@
The goal of Section 4 is to demonstrate the remarkable properties of the Lasso and the Thresholded Lasso estimators: while exact recovery of all non-zero variables requires very stringent incoherence and $\beta_{\min}$ conditions, we can significantly relax both conditions when we only require a subset $A_{0}$ of active variables to be included in our selection set. We loosely refer to $A_{0}$ or its superset $T_{0}\supseteq A_{0}$ which we aim to identify as an active set throughout this work. When $\beta_{\min,A_{0}}$ (15) is sufficiently large, we have $A_{0}\subset I$ while achieving the sparse oracle inequalities in the sense of (11); cf. (4.3) and Theorem 4.4.
One of the reviewers brought to our attention that [43] provide bounds similar to the Dantzig selector under the upper and lower sparse Riesz condition (SRC), which are similar to the upper and lower bounds in (2.6) in the present work. Both models allow potentially many small coefficients in the true $\beta$. Specifically, we design a new set of experiments to evaluate the impact of sparsity $s$ and the $\beta_{\min,A_{0}}$ (e.g., $C_{a}\lambda\sigma$ in (72)) condition on the recovery of the first $s_{0}$ components in $\beta$, where we set $\beta_{T_{0}^{c}}$ (with support size $s-s_{0}$) to have a fixed $\ell_{2}$ norm but potentially many small coordinates with magnitude $\asymp\sigma/\sqrt{n}$. See Section 5 for the setup of numerical simulations.
Not included in the present work are Theorem 2.1 and the Iterative Procedure by [47], where we show conditions under which one can recover a sparse subset of strong signals when $\beta_{\min}:=\min_{j\in S}\left\lvert\beta_{j}\right\rvert\geq C\sigma\sqrt{{2s\log p}/{n}}$, where $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$, $s=\left\lvert S\right\rvert$ and $C$ depends on the restricted eigenvalue parameter; cf. Theorems 3.1 in [48]. When $\beta_{\min}$ is sufficiently large, the range of thresholding parameters is even more flexible, which we elaborate in [47] and [48], cf. Theorem 3.1, and hence details are omitted from the current paper. We do show numerical examples for which the Thresholded Lasso recovers the support $S$ exactly with high probability, using a small number of samples per non-zero component in $\beta$, for which the Lasso would certainly have failed, as predicted by the work of [38, 37]. These result have been presented in part in an earlier version of the present paper [48] and the conference paper by [47], which was inspired by an oracle result on the $\ell_{2}$ loss for the Dantzig selector by [9]; cf. the supplementary Proposition I.4.
@@
Organization of the paper. The rest of the paper is organized as follows. In Section 2, we describe a thresholding framework for the general setting, and highlight the role thresholding plays in terms of recovering the best subset of variables; we present the main Theorem 2.1 and oracle results for the Lasso estimator, which are crucial in proving Theorem 2.1. We discuss background and related work in Section 2.3. Section 3 provides the proof sketch of the main Theorem 2.1 and the proof of Theorem 2.1 appears in Section A.1. Section 4 discusses Type II errors and the $\ell_{1}$ and $\ell_{2}$ loss. Section 5 and the supplementary Section K include simulation results. We prove Lemmas 2.5 and 2.7 in Section B. We conclude in Section 6. Additional technical proofs are included in the supplement.

### v3

@@ 1 Introduction
Here $X$ is an $n\times p$ design matrix, $Y$ is a vector of noisy observations, and $\epsilon$ is the noise term. We assume throughout this paper that $p\geq n$ (i.e. high-dimensional) and $\epsilon$ is a vector of i.i.d. normal $N(0,\sigma^{2})$ random variables. Denote by $p]=\{1,\ldots,p\}$. The notation $\left\lVert\beta\right\rVert_{2}=(\sum_{j}\beta_{j}^{2})^{1/2}$ stands for the $\ell_{2}$ norm of $\beta$. Given such a linear model, two key tasks are: (1) to select the relevant set of variables and (2) to estimate $\beta$ with bounded $\ell_{2}$ loss. In particular, recovery of the sparsity pattern $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta):=\left\{j\,:\,\beta_{j}\neq 0\right\}$, also known as variable (model) selection, refers to the task of correctly identifying the support set, or a subset of “significant” coefficients in $\beta$, based on the noisy observations. Here and in the sequel, we assume that each column vector $X_{j}\in\mathbb{R}^{n},j\in[p]$ of the fixed design matrix $X$ has $\ell_{2}$-length $\left\lVert X_{j}\right\rVert_{2}=\sqrt{n}$.
Even in the noiseless case, recovering $\beta$ (or its support) from $(X,Y)$ seems impossible when $n\ll p$ given that we have more variables than observations. Over the past two decades, a line of research shows that when $\beta$ is sparse, that is, when it has a relatively small number of nonzero coefficients, and when design matrix $X$ behaves sufficiently nicely in a sense that it satisfies certain incoherence conditions, it becomes possible to reconstruct $\beta$ [12, 8, 9].
@@
where $\upsilon_{J}$ represents the subvector of $\upsilon\in\mathbb{R}^{p}$ confined to a subset $J$ of $[p]$. It is clear that as $k_{0}$ and $s_{0}$ become smaller, this condition is easier to satisfy. To be clear, RE conditions alone are not sufficient for the Lasso to recover the model $S$ exactly. Moreover, to ensure variable selection consistency, a
$\text{ $\beta_{\min}$ condition:}\quad\min_{j\in S}\left\lvert\beta_{j}\right\rvert\geq C\sigma\sqrt{{2\log p}/{n}},$ (4)
@@
    Threshold the initial Lasso estimator $\beta_{\text{\rm init}}$ with $t_{0}$, with the general goal such that, we get a set $I$ with cardinality at most $2s$; in general, we also have $|I\cup S|\leq 2s$, where $I=\left\{j\in[p]:\left\lvert\beta_{j,\text{\rm init}}\right\rvert\geq t_{0}\right\}$ for some $t_{0}\asymp\sigma\sqrt{2\log p/n}$ with hidden constant to be specified;
@@
In Theorem 2.1, we show that the critical threshold level for estimating a high dimensional sparse vector $\beta\in\mathbb{R}^{p}$ should be set at the level $t_{0}\asymp\lambda\sigma$ for $\lambda=\sqrt{2\log p/n}$, to retain signals in $\beta_{\text{\rm init}}$ at or above the level of $\lambda\sigma$, where $\beta_{\text{\rm init}}$ is the solution to the Lasso estimator (2) obtained in Step 1. Moreover, we show that $d_{0}$ and $t_{0}$ are allowed to depend on sparse and restricted eigenvalue parameters of the design matrix; cf. Section 3.1. From a theoretical point of view, the framework for our analysis is set upon the Restricted Eigenvalue type of condition and an upper sparse eigenvalue (USE) condition, namely,
@@
$\text{(USE)}\quad\Lambda_{\max}(2s)\;\stackrel{{\scriptstyle\triangle}}{{=}}\;\max_{\upsilon\not=0;2s-\text{sparse}}\;\left\lVert X\upsilon\right\rVert_{2}^{2}/(n\left\lVert\upsilon\right\rVert_{2}^{2})<\infty.$ (5)
@@
These are among the most general assumptions on the design matrix, guaranteeing sparse recovery and oracle inequalities in the $\ell_{2}$ loss for the Lasso estimator as well as the Dantzig selector: $\textnormal{{RE}}(s,k_{0},X)$ is shown to be a relaxation of the restricted isometry property (RIP) under suitable choices of parameters involved in each condition [9, 3]. Part of this work was presented in a conference paper by [47]. Importantly, we present significant and novel extensions in both theory and numerical simulations, with regards to the Thresholded Lasso and the Lasso under the RE and sparse eigenvalue conditions.
@@
While the crucial theoretical and methodological ideas presented here originate from [47], the current work significantly expands the original ideas and show new results on the sparse oracle inequalities in Theorems 2.1 and 2.4. Compared with the original paper [47], we further study the behavior of the Thresholded Lasso in several challenging situations in Sections 45, and the supplementary Section I. We find that our estimator is robust and adaptive to the overwhelming presence of weak signals in $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ in both theoretical and practical senses. We show that the Thresholded Lasso tradeoffs false positives and false negatives nicely in this case: its advantage in terms of model selection over the Lasso and adaptive Lasso [52, 51] is clearly evident by examining their ROC curves empirically. Our numerical simulations in Section I show that the rates for exact recovery of the support rise sharply for a few types of random matrices once the number of samples passes a certain threshold, using the Thresholded Lasso estimator.
Notation and definitions. Let $T\subseteq[p]$ be a fixed subset of indices. Let $X_{T}$ be the $n\times\left\lvert T\right\rvert$ submatrix obtained by extracting columns of $X$ indexed by $T$. We use $\upsilon_{T}$ to represent the subvector of $\upsilon\in\mathbb{R}^{p}$ confined to a subset $T$ of $[p]$. Let $\lVert\upsilon\rVert_{2}^{2}=\sum_{j=1}^{p}\upsilon_{j}^{2}$. Occasionally, we use $\upsilon_{T}\in\mathbb{R}^{\left\lvert T\right\rvert}$ to also represent its 0-extended version $\upsilon^{\prime}\in\mathbb{R}^{p}$ such that $\upsilon^{\prime}_{T^{c}}=0$ and $\upsilon^{\prime}_{T}=\upsilon_{T}$. We will use $\beta^{\mathop{\natural}\kern 0.60275pt}(T)\in\mathbb{R}^{p}$ to represent the 0-extended version of $\beta_{T}\in\mathbb{R}^{\left\lvert T\right\rvert}$ such that $\beta_{T}^{\mathop{\natural}\kern 0.60275pt}({T})=\beta_{T}$ and $\beta_{T^{c}}^{\mathop{\natural}\kern 0.60275pt}({T})=0$. For a matrix $A$, let $\Lambda_{\min}(A)$ and $\Lambda_{\max}(A)$ denote the smallest and the largest eigenvalues respectively. Let $s=|S|$. We assume the lower sparse eigenvalue (LSE) condition, namely,
@@
$\text{(LSE)}\quad\Lambda_{\min}(2s)\;\stackrel{{\scriptstyle\triangle}}{{=}}\;\min_{\upsilon\not=0;2s-\text{sparse}}\;\;\left\lVert X\upsilon\right\rVert_{2}^{2}/(n\left\lVert\upsilon\right\rVert_{2}^{2})>0,$ (6)
@@
${\left\lvert{\;\langle{\,X_{T}\upsilon,X_{T^{\prime}}{\upsilon^{\prime}}\,}\rangle\;}\right\rvert}/{n}\leq\theta_{s,s^{\prime}}\left\lVert\upsilon\right\rVert_{2}\left\lVert\upsilon^{\prime}\right\rVert_{2},\;\text{ where }s+s^{\prime}\leq p,$ (7)
$\text{ and }\;\theta_{s,s^{\prime}}\leq(\Lambda_{\max}(s)\Lambda_{\max}(s^{\prime}))^{1/2}\;\text{ by the Cauchy Schwarz inequality.}$ (8)
Note that small values of $\theta_{s,s^{\prime}}$ indicate that disjoint subsets of covariates in $X_{T}$ and $X_{T^{\prime}}$ span nearly orthogonal subspaces. Moreover, we have $\theta_{s,s^{\prime}}\leq{(\Lambda_{\max}(s+s^{\prime})-\Lambda_{\min}(s+s^{\prime}))}/{2}$; cf. Lemma 2.8. Technically speaking, each of the entities defined above, namely, $1/\Lambda_{\min}(2s)$, $\Lambda_{\max}(2s)$, $\theta_{s,s^{\prime}}$, and $K(s,k_{0})$ as introduced in (3), is a non-decreasing function of $s$, $s^{\prime}$, and $k_{0}$. Nonetheless, we crudely consider these as constants following how they are typically treated in the literature as it is to be understood that they grow very slowly with $s$ and $s^{\prime}$; see for example [8, 9][26], and [3]. For two numbers $a,b$, $a\wedge b:=\min(a,b)$, and $a\vee b:=\max(a,b)$. We write $a\asymp b$ if $ca\leq b\leq Ca$ for some positive absolute constants $c,C$ which are independent of $n,p$, and $\gamma$. We write $f=O(h)$ or $f\ll h$ if $\left\lvert f\right\rvert\leq Ch$ for some absolute constant $C<\infty$ and $f=\Omega(h)$ or $f\gg h$ if $h=O(f)$. We write $f=o(h)$ if $f/h\to 0$ as $n\to\infty$.
@@ 1.1 Sparse oracle inequalities
[…]
Formally, we evaluate the selection set through the following criterion. Consider the least squares estimators $\hat{\beta}_{I}=(X_{I}^{T}X_{I})^{-1}X_{I}^{T}Y$, where $I\subset[p]$ and |I|s|I|\ $\leq$ s. Here and in the sequel, let $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)=\hat{\beta}_{I}$ and $\hat{\beta}_{I^{c}}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)=0$. Consider the ideal least-squares estimator $\beta^{\diamond}$ based on a subset $I$ of size at most $s$, which minimizes the mean squared error:
@@
${\textbf{E}}\left\lVert\beta-\beta^{\diamond}\right\rVert_{2}^{2}$ $\geq$ $\min\left(1,{1}/{\Lambda_{\max}(s)}\right)\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n),\text{ where }$ (10)
$\sum_{i=1}^{p}\min(\beta_{i}^{2},\sigma^{2}/n)$ $=$ $\min_{I\subseteq[p]}\left\lVert\beta-\beta^{\mathop{\natural}\kern 0.60275pt}({I})\right\rVert_{2}^{2}+{|I|\sigma^{2}}/{n}$
@@
in view of (10); cf. the supplementary Section A. Here the $\ell_{2}$ loss in (11) is optimal up to a $\log p$ factor. We note that (12) is not the tightest upper bound that we could derive due to a relaxation we have on the lower bound as stated in (10). Nevertheless, we use it for its simplicity.
Essential sparsity and objectives. The current paper answers the following question: Is there a good thresholding rule that enables us to obtain a sufficiently sparse estimator $\hat{\beta}$ that satisfies an oracle inequality in the sense of (11), when some components of $\beta_{S}$ are well below $\sigma/\sqrt{n}$? Such oracle results are accomplished without any knowledge of the significant coordinates or parameter values of $\beta$. Both Theorem 2.1 and the supplementary Theorem G.3 answer this question positively, where we elaborate upon the sparse recovery properties of the Lasso and Dantzig selector in combination with thresholding and refitting.
For a given pair of $(n,p)$ values, the essential sparsity parameter $s_{0}$ characterizes more accurately than $s$, the number of significant coefficients of $\beta$ with respect to the noise level $\sigma$ that we should (could) try to recover. Denote by $s_{0}$ the smallest integer such that the following holds [9]:
@@
The parameter $s_{0}$ is relevant especially when we do not wish to impose any lower bound on $\beta_{\min}$. This is the focus of Theorem 2.1. To make this statement precise, we have as a consequence of the definition in (13),
@@
The Thresholded Lasso algorithm is constructive in that it relies neither on the unknown parameters $\left\lvert S\right\rvert$ or $\beta_{\min}:=\min_{j\in S}|\beta_{j}|$, nor the exact knowledge of those that characterize the incoherence conditions on $X$. Instead, our choices of penalty $\lambda_{n}$ (Step 1) and threshold $t_{0}$ (Step 2) only depend on $\sigma,n$, and $p$, and some crude estimation of certain sparse eigenvalue parameters; cf. Section 3.1. In practical settings, one can choose $\lambda_{n}$ using cross-validation; See for example the subsequent work by [50], where we use cross-validation to choose both penalty and threshold parameters in the context of covariance selection based on Gaussian graphical models.
In Section 4, we briefly discuss possibilities of recovering a subset of strong signals via thresholding, despite the existence of (or influence from) other relatively weaker signals. Let $T_{0}$ denote the largest $s_{0}$ coordinates of $\beta$ in absolute values. As a consequence of the definition in (13), we have $T_{0}=\{1,\ldots,s_{0}\}$ and $|\beta_{j}|<\lambda\sigma$ for all $j\in T_{0}^{c}$; cf. (14). More precisely, we decompose $T_{0}=\{1,\ldots,s_{0}\}$ into two sets: $A_{0}$ and $T_{0}\setminus A_{0}$, where $A_{0}$ contains the set of coefficients of $\beta$ strictly above $\lambda\sigma$, for which we define a constant $\beta_{\min,A_{0}}$:
@@
The goal is to demonstrate the remarkable properties of the Lasso and the Thresholded Lasso estimators: while exact recovery of all non-zero variables requires very stringent incoherence and $\beta_{\min}$ conditions, we can significantly relax both conditions when we only require a subset $A_{0}$ of active variables to be included in our selection set. We loosely refer to $A_{0}$ or its superset $T_{0}\supseteq A_{0}$ which we aim to identify as an active set throughout this work. When $\beta_{\min,A_{0}}$ (15) is sufficiently large, we have $A_{0}\subset I$ while achieving the sparse oracle inequalities in the sense of (11); cf. (4.3) and Theorem 4.4.
One of the reviewers brought to our attention that [43] provide bounds similar to the Dantzig selector under the upper and lower sparse Riesz condition (SRC), which are similar to the upper and lower bounds in (31) in the present work. Both models allow potentially many small coefficients in the true $\beta$. Specifically, we design a new set of experiments to evaluate the impact of sparsity $s$ and the $\beta_{\min,A_{0}}$ (e.g., $C_{a}\lambda\sigma$ in (69)) condition on the recovery of the first $s_{0}$ components in $\beta$, where we set $\beta_{T_{0}^{c}}$ (with support size $s-s_{0}$) to have a fixed $\ell_{2}$ norm but potentially many small coordinates with magnitude $\asymp\sigma/\sqrt{n}$. See Section 5 for the setup of numerical simulations.
Not included in the present work are Theorem 2.1 and the Iterative Procedure by [47], where we show conditions under which one can recover a sparse subset of strong signals when $\beta_{\min}:=\min_{j\in S}\left\lvert\beta_{j}\right\rvert\geq C\sigma\sqrt{{2s\log p}/{n}}$, where $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$, $s=\left\lvert S\right\rvert$ and $C$ depends on the restricted eigenvalue parameter; cf. Theorems 3.1 in [48]. When $\beta_{\min}$ is sufficiently large, the range of thresholding parameters is even more flexible, which we elaborate in [47] and [48], cf. Theorem 3.1, and hence details are omitted from the current paper. We do show numerical examples for which the Thresholded Lasso recovers the support $S$ exactly with high probability, using a small number of samples per non-zero component in $\beta$, for which the Lasso would certainly have failed, as predicted by the work of [38, 37]. These results have been presented in part in an earlier version of the present paper [48] and the conference paper by [47].
@@
Organization of the paper. The rest of the paper is organized as follows. In Section 2, we describe a thresholding framework for the general setting, and highlight the role thresholding plays in terms of recovering the best subset of variables; we present the main Theorem 2.1 and oracle results for the Lasso estimator. We discuss related work in Section 2.3. Section 3 provides the proof sketch of Theorems 2.1 and 2.4, and the proof of Theorem 2.1 appears in Section 6.1. Section 4 discusses Type II errors and the $\ell_{1}$- and $\ell_{2}$-loss. Section 5 and the supplementary Section I include simulation results. We prove Lemmas 2.52.74.2, and Lemma 4.3 in Sections 6.4 to 6.7. We give a sketch proof of Theorem 3 [43] in the supplementary Section 7. We conclude in Section 8. Additional technical proofs are included in the supplement.